OpenStudy (experimentx):

Integrate:- \[ \int_1^\infty \frac{1}{\sqrt{x(x^2-1)}}dx \]

hartnn (hartnn):

will try 1) x = 1/y 2) Beta Function

OpenStudy (experimentx):

looks like it might work ... should have realized when i saw fraction of gamma functions.

hartnn (hartnn):

infact x^2 = 1/y

OpenStudy (experimentx):

same thing ... you can make replacements later.

hartnn (hartnn):

yup, now we just need to find the exponents of y and 1-y

OpenStudy (experimentx):

since wolf gave me elliptic functions as results of indefinite results it was bugging me.

hartnn (hartnn):

1/ sqrt (x (x-1)) would be more interesting.....it won't converge...

OpenStudy (experimentx):

it is same as 1/x ... whose integral is log(x) ... as x->inf ... log(x) diverges.

hartnn (hartnn):

:D

OpenStudy (experimentx):

(y)